OpenStudy (anonymous):

What is the value of x, when 10(x + 2) = 5(x + 8)? A) x = -2 B) x = 2 C) x = 4 D) x = 5/6

OpenStudy (mathhelp346):

10(x+2) = 5(x+8) 10x+20 = 5x+40 10x = 5x+20 5x = 20 x = 4